Consider a solution that is 1 0 times 10^-2 M in B
Answer #1

BaSO4 Ksp = 1.1*10^-10

CaSO4 KSp = 2.4*10^-5

first to precipitate is BaSO4

so

Ksp = [Ba+2][SO4-2]

1.1*10^-10 = (10^-2)(M)

M =(1.1*10^-10)/(10^-2) = 1.1*10^-8 M of Na2SO4